OpenStudy (anonymous):

how do i solve 0.65x+0.3x=x-3?

OpenStudy (anonymous):

am i supose to combine like terms like 0.95x=x-3?

OpenStudy (jannine24):

yes first you combine like terms

OpenStudy (jannine24):

0.95x = x-3

OpenStudy (jannine24):

then what do you do

OpenStudy (anonymous):

get x on the right side to go to left side?

OpenStudy (jannine24):

yes

OpenStudy (anonymous):

+x on the right side to make it 1.95x=3?

OpenStudy (jannine24):

so you would get -1x + 0.95x =-3

OpenStudy (jannine24):

then subtract -1 from .095 you get -.05

OpenStudy (jannine24):

so now you have -0.05x=-3

OpenStudy (jannine24):

what do you do now

OpenStudy (anonymous):

divide both sides by 0.5?

OpenStudy (jannine24):

but -0.05

OpenStudy (anonymous):

will it be a positive since both are negative?

OpenStudy (jannine24):

yes it is a positive

OpenStudy (anonymous):

x=6?

OpenStudy (jannine24):

not 60 you get 60

OpenStudy (jannine24):

then go back and check your answer to see if you are right 0.65(60) + 0.3(60)= 60-3

OpenStudy (jannine24):

so you get 39+18=57

OpenStudy (anonymous):

ok i see it now thank you for steps

OpenStudy (jannine24):

no problem